A 6-footer Pedro lifted a 50-kg sack of rice and put it on top of his head. How much power does Pedro have if he was able to lift the sack in 1 second?


a. 1000 J/s

b. 9000 J/s

c. 100 J/s

d. 900 J/s


1
Expert's answer
2022-03-09T10:33:15-0500

The power

"P=\frac{W}{t}=\frac{mgh}{t}\\
=\rm \frac{50\: kg*9.8\: N/kg*1.8\: m}{1\: s}=882\: J/s=900\: J/s"

Answer: d. 900 J/s